# Language-Agnostic Modeling of Source Reliability on Wikipedia

### v2

@@ Abstract.
Over the last few years, content verification through reliable sources has become a fundamental need to combat disinformation. Here, we present a language-agnostic model designed to assess the reliability of sources across multiple language editions of Wikipedia. Utilizing editorial activity data, the model evaluates source reliability within different articles of varying controversiality such as Climate Change, COVID-19, History, Media, and Biology topics. Crafting features that express domain usage across articles, the model effectively predicts source reliability, achieving an F1 Macro score of approximately 0.80 for English and other high-resource languages. For mid-resource languages, we achieve 0.65 while the performance of low-resource languages varies; in all cases, the time the domain remains present in the articles (which we dub as permanence) is one of the most predictive features. We highlight the challenge of maintaining consistent model performance across languages of varying resource levels and demonstrate that adapting models from higher-resource languages can improve performance. This work contributes not only to Wikipedia’s efforts in ensuring content verifiability but in ensuring reliability across diverse user-generated content in various language communities.
@@
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Wikisccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
The proliferation of fake news in recent years is deteriorating the reliability and trustworthiness of the online information ecosystem (Del Vicario et al., 2016; Vosoughi et al., 2018). To mitigate their spread and impact in society, researchers on web and data mining are devoting a great deal of effort in generating resources to detect untrue content on social media platforms (Su et al., 2023; Xu et al., 2022; Ashkinaze et al., 2024; Xiao et al., 2024; Guo et al., 2023b; Xiao et al., 2023; Wu et al., 2023).
In popular web platforms like Wikipedia [^1] Wikipedia is among the top visited websites worldwide in 2024: https://www.semrush.com/trending-websites/global/all., disinformation is also a main threat to knowledge integrity (Aragón and Sáez-Trumper, 2021). Wikipedia’s conception of knowledge as a service requires content to be verifiable in order to preserve its reliability and integrity [^2] https://research.wikimedia.org/knowledge-integrity.html. Thus, the platform’s policy on verifiability implies that readers should be able to check that all the information in Wikipedia articles comes from reliable sources [^3] https://en.wikipedia.org/wiki/Wikipedia:Verifiability.. While a classical patrolling technique in this platform is to detect and remove statements that violate basic core content policies (Kumar et al., 2016), some Wikipedia editors have proven more effective in tackling disinformation by first identifying unreliable sources (Cohen, 2021). Recent research revealed the positive impact of the community-curated list of perennial sources in English Wikipedia (Baigutanova et al., 2023a). However, that list is limited or even non-existent in other language editions (Baigutanova et al., 2023b).
To address this challenge, we present our research on source reliability, based on language-agnostic approaches and edit activity data from multiple Wikipedia language editions. We aim to support Wikipedia editors, including the ones in medium and small projects that often miss advanced tools, in identifying sources with patterns associated with low reliability, and in monitoring their prevalence across language editions. To meet this goal, we address the following research questions:
@@ 2. Related Work
Several research works have assessed the verifiability of Wikipedia content and the reliability of the sources by analysing, e.g., the quality of references, the appropriateness of links between articles, and the consensus between editors regarding this content.
@@
The Contropedia project (Borra et al., 2015), for example, aimed at identifying and visualizing controversial links between Wikipedia articles, i.e., links that point to other articles in the platform and whose collaborative edit history shows signs of dispute between editors before reaching consensus. Furthermore, several tools provide interactive visualizations that help understand the path to consensus for a Wikipedia article  (Borra et al., 2015; Guo et al., 2023a). Cultural differences in the controversiality of specific topics have been shown to exist (e.g., (Fu et al., 2023)).
The Citation Detective tool (Chou et al., 2020) leveraged machine learning models (Redi et al., 2019) to assess the evolution of reference quality between 2010 and 2020, in terms of the ratio of sentences missing a required citation, and the ratio of citations that reference non-authoritative sources. The authors found out that, thanks to the efforts of the community, the ratio of sentences missing a required citation was significantly reduced, while the ratio of references to non-authoritative sources has remained below $1\%$. More recently,  (Petroni et al., 2023) has proposed a neural-network-based system to improve the verifiability of content.
The observed improvements in reference quality coincide with the introduction of the perennial source list (Wikipedia contributors, 2024), a curated, non-exhaustive list of sources whose reliability and use on Wikipedia are frequently discussed. Though the perennial source list has been used by several works to assess the authoritativeness of sources, only a few language versions of Wikipedia contain lists of perennial sources and, as the reliability of a source is not static, these lists might evolve through time. Moreover, (Baigutanova et al., 2023b) showed that the reliability of references might be different across cultures, such that unreliable references in one language might persist across editions in another language. This makes it difficult to train general models for source controversiality and, given that some language editions of Wikipedia are still small, achieving reference quality across all languages calls for language-agnostic approaches as the one presented in this work. Beyond the perennial source list, other domain quality rankings for news sources exist, such as the Iffy index of unreliable news sources  (Barret, 2021) based on the Media Bias/Fact-Check (Zandt, 2024). However, we show that the latter corresponds only loosely with the perennial source list, pointing to the necessity of Wikipedia-specific models.
[…]
@@
The multilingual nature of Wikipedia calls for specific modeling techniques. Models such as multilingual BERT have been used to detect vandalism (Trokhymovych and Saez-Trumper, 2021) or political bias (Swati et al., 2023), while other works have built rankings of relative quality in multiple language versions of the platform (Lewoniewski et al., 2020, 2019). For some other tasks, and due to the limitations found in languages with smaller Wikipedia datasets, language-agnostic approaches have been used, e.g., to assess article quality (Das et al., 2024), or to perform automatic entity-linking (Gerlach et al., 2021) and topic classification (Johnson et al., 2021). Language-agnostic models using features extracted from the network structure or user behavior have also been used in the more general context of detecting misinformation or low-quality information in social media. In this line, Shu et al. (2019) constructed embeddings of users and news and used diffusion models to trace the paths through which fake news propagates, while Zhao et al. (2021) extracted features from the user interaction network to help identify misinformation in online health communities. We show that, indeed, user-behavior related language-agnostic signals are useful for modeling source reliability, and that they can be adapted across languages to improve performance in low-resource settings.
@@ 3.1. Perennial Sources
We use the English Wikipedia perennial source list (Wikipedia contributors, 2024) as ground truth for modeling source reliability. It was collected on March 30, 2023. Although the list exists for 12 other languages, we only use the English version as it is by far the most extensive (see (Baigutanova et al., 2023b) for details on coverage). In English, it consists of a collection of web domains in five categories: blacklisted (569), deprecated (125), generally unreliable (178), no consensus (137), and generally reliable (184). To build a binary classifier, we use the generally reliable domains as positive labels, merge the domains blacklisted, deprecated, and generally unreliable into a single unreliable category (i.e., the negative label) and ignore the no consensus domains.
Table 1. Statistics of the topic-language datasets used in the study grouped by high, mid, and low resource languages. All numbers are averages (apart from the number of languages). Perennial Domains refers to the number of domains for the perennial source list in English found in the different datasets.
Topic # Langs # Articles Revs Revs/Article URLs URLs/Article Domains Perennial Domains
High Climate Change 7 $1378$ 372 133372133372\,133372 133 $269$ 58 6715867158\,67158 671 $49$ 15 3631536315\,36315 363 $203$
COVID-19 7 $898$ 233 905233905233\,905233 905 $260$ 60 1686016860\,16860 168 $89$ $8289$ $206$
Biology Sample 7 13 6071360713\,60713 607 586 171586171586\,171586 171 $43$ 59 7635976359\,76359 763 $6$ 11 7131171311\,71311 713 $148$
History Sample 7 $5710$ 624 428624428624\,428624 428 $109$ 52 7025270252\,70252 702 $12$ 12 1211212112\,12112 121 $166$
Media Sample 7 $5577$ 819 541819541819\,541819 541 $146$ 108 583108583108\,583108 583 $22$ 21 2902129021\,29021 290 $259$
Mid Climate Change 36 $395$ 28 3642836428\,36428 364 $71$ $6224$ $19$ $2599$ $94$
COVID-19 37 $272$ 18 5231852318\,52318 523 $68$ $8984$ $39$ $1788$ $102$
Biology 37 $4153$ 83 5998359983\,59983 599 $20$ $8691$ $3$ $2022$ $61$
History 37 $1680$ 83 2588325883\,25883 258 $49$ $6274$ $7$ $2054$ $67$
Media 37 $1018$ 46 9574695746\,95746 957 $46$ $7932$ $10$ $2416$ $123$
@@
Biology 41 $593$ 21 0962109621\,09621 096 $35$ $1159$ $4$ $460$ $16$
History 55 $511$ 13 5011350113\,50113 501 $26$ $952$ $6$ $406$ $22$
Media 70 $159$ $2262$ $14$ $528$ $7$ $264$ $30$
@@ 3.2. Language Groups
As we aim to test the proposed model on a variety of languages, we collect data on all $326$ available language editions of Wikipedia. We then discard all languages which have fewer then two reliable and unreliable sources according to the English perennial source list. Among the remaining ones, we distinguish between low, middle, and high-resource languages by examining the number of “active users”, defined as “registered users who have made at least one edit in the last $30$ days.” [^4] https://meta.wikimedia.org/wiki/List_of_Wikipedias (accessed on 01-03-2024) We consider the top 5% of the languages having the most active users as high-resource, the subsequent 25% as mid-resource, and the remaining 70% as low-resource languages.
@@ 3.3.1. Topic Selection
In this study, we have focused on articles from the following five topic datasets: Climate Change, COVID-19, Biology, History, and Media. Climate Change and COVID-19 were selected as they are both controversial but for a different time duration. COVID-19 appeared at the end of 2019, while Climate Change has been a controversial topic for decades (Korte et al., 2023). Both have a Wikiproject community of editors  [^5] https://en.wikipedia.org/wiki/Wikipedia:WikiProject who label articles as relevant to the topic with special guidelines. For comparison, we also selected three topics that are not as controversial, but which varied in terms of reference quality based on the “reference risk” metric (which is also based on the usage of perennial sources) from previous work (Baigutanova et al., 2023a). We choose one example from each risk category: Biology (low risk), History (medium risk), and Media (high risk).
@@ 3.3.2. Data Download
For the first two topics, we use the articles listed by the corresponding topic-based WikiProjects. For the latter three, we use the topics identified using the Wikipedia Objective Revision Evaluation Service (ORES) machine learning system (Halfaker and Geiger, 2020). To keep the sizes of datasets comparable and manageable, we sample the articles identified by ORES to be in Biology and History topics by 10% (each), and in the Media topic by 2%. These two approaches allow us to find articles in the English Wikipedia for each of the five topics, resulting in a total of 84 0278402784\,02784 027 articles. To find the versions of these articles in other languages, we use the MediaWiki API, which results in another 335 809335809335\,809335 809 of articles in all of the 326 languages. We collect the articles and their complete edit history using the MediaWiki API in between December 2023 and February 2024.
@@ 3.3.3. Extraction of Source Domains
Citations may appear in the body of the article text and as references at the bottom of the article, each of which has standard templates. Each revision that involves a citation contains the author of the revision, position within the document of the citation, text of the citation, and meta information, such as whether it is a DOI, URL, ISBN, ISSN, etc. or a raw URL. We standardize these revisions to create a list of “source edits” containing the metadata about each insertion, deletion, or edit that involves a source. We resolve redirects, apply rules to clean and standardize the URLs (including DOI references redirecting to a resource) and merge revisions when the same user edits an article in succession (regardless of the time). Finally, we extract the domains from the URLs, obtaining 376 566376566376\,566376 566 unique source domains.
@@ 3.3.4. Dataset Statistics
The summary statistics about the topic-language datasets (or simply “datasets” throughout this paper), for the five topics grouped into low-, mid- and high-resource languages can be seen in Table 1. Recall that we only use datasets that have passed the threshold of having at least 2 reliable and 2 unreliable perennial sources.
@@ 3.4. Feature Definition
Unlike previous studies measuring Wikipedia edit quality that did not consider certain types of edits and reverts (Trokhymovych et al., 2023), we propose to focus on the entire edit history concerning each domain. Using edit metadata, we define 52 features capturing the popularity of the domain, the “permanence” of the domain across edits, and the number and type of users involved in edits adding or removing that domain. As the features are computed for each domain within each dataset separately (recall, by “dataset” we mean a collection of articles on a certain topic, in a certain language) of varying sizes and editorial frequencies, we explore different ways of normalizing these statistics w.r.t. the first appearance of the domain, age of the dataset, and in terms of time duration or number of revisions.

### v3

@@ Abstract.
Over the last few years, verifying the credibility of information sources has become a fundamental need to combat disinformation. Here, we present a language-agnostic model designed to assess the reliability of web domains as sources in references across multiple language editions of Wikipedia. Utilizing editing activity data, the model evaluates domain reliability within different articles of varying controversiality, such as Climate Change, COVID-19, History, Media, and Biology topics. Crafting features that express domain usage across articles, the model effectively predicts domain reliability, achieving an F1 Macro score of approximately 0.80 for English and other high-resource languages. For mid-resource languages, we achieve 0.65, while the performance of low-resource languages varies. In all cases, the time the domain remains present in the articles (which we dub as permanence) is one of the most predictive features. We highlight the challenge of maintaining consistent model performance across languages of varying resource levels and demonstrate that adapting models from higher-resource languages can improve performance. We believe these findings can assist Wikipedia editors in their ongoing efforts to verify citations and may offer useful insights for other user-generated content communities.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Wikisccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
The proliferation of fake news in recent years is deteriorating the reliability and trustworthiness of the online information ecosystem (Del Vicario et al., 2016; Vosoughi et al., 2018). To mitigate their spread and impact in society, researchers in web and data mining are devoting a great deal of effort to generating resources to detect untrue content on social media platforms (Su et al., 2023; Xu et al., 2022; Ashkinaze et al., 2024; Xiao et al., 2024; Guo et al., 2023b; Xiao et al., 2023; Wu et al., 2023).
In popular web platforms like Wikipedia [^1] Wikipedia is among the top visited websites worldwide: https://www.semrush.com/trending-websites/global/all., disinformation is also a main threat to knowledge integrity (Aragón and Sáez-Trumper, 2021). Wikipedia’s conception of knowledge as a service requires content to be verifiable to preserve its reliability and integrity [^2] https://research.wikimedia.org/knowledge-integrity.html. Thus, the platform’s policy on verifiability implies that readers should be able to check that all the information in Wikipedia articles comes from reliable sources [^3] https://en.wikipedia.org/wiki/Wikipedia:Verifiability.. While a classical patrolling technique in this platform is to detect and remove statements that violate basic core content policies (Kumar et al., 2016), some Wikipedia editors have proven more effective in tackling disinformation by first identifying unreliable sources (Cohen, 2021). Recent research revealed the positive impact of the community-curated list of perennial sources in English Wikipedia (Baigutanova et al., 2023a). However, that list is limited or even non-existent in other language editions (Baigutanova et al., 2023b). A great deal of manual effort would be necessary to create such community-curated resources in all of the languages on Wikipedia.
To address this challenge, we propose a language-agnostic approach to reveal source reliability that leverages the implicit signals in the edit activity data from multiple Wikipedia language editions. We aim to support Wikipedia editors, including those in medium and small projects that often miss advanced tools, in identifying sources with patterns associated with low reliability, and in monitoring their prevalence across language editions. To meet this goal, we address the following research questions:
@@ 2. Related Work
This work approaches the challenge of measuring source reliability on Wikipedia through the lens of controversiality metrics. This approach builds on the methodology of the Contropedia project (Borra et al., 2015), aimed at identifying and visualizing controversial links between Wikipedia articles, i.e., links that point to other articles in the platform and whose collaborative edit history shows signs of dispute between editors before reaching consensus. This is not the only effort in this space. Systems like Edit-History Vis (Guo et al., 2023a) have also included interactive visualizations to illustrate the path to consensus for a Wikipedia article. Cultural differences in the controversiality of specific topics have been shown to exist (e.g., (Rogers et al., 2012; Fu et al., 2023)). Unlike these previous approaches primarily focusing on wikilinks within article text, our work shifts attention to citations, aiming to explore the reliability of sources in Wikipedia.
Previous research has examined assessed the verifiability of Wikipedia content (Chen and Roth, 2012; Benjakob et al., 2022; Halitaj and Zubiaga, 2024, 2025) and the reliability of the sources (Lewoniewski et al., 2020; Yang and Colavizza, 2024).  Baigutanova et al. (2023a) addressed both goals by analyzing reference quality on English Wikipedia over time using two key metrics: the proportion of sentences lacking necessary citations, as identified by machine learning models (Redi et al., 2019; Chou et al., 2020), and the proportion of citations linking to non-authoritative sources. For the latter, non-authoritative sources were determined using the perennial source list (Wikipedia contributors, 2024), a curated list of sources whose reliability and use on Wikipedia are frequently discussed. The study found the share of references to non-authoritative sources has remained below 1%, with a notable decline following the introduction of the list in 2018.
Although the perennial source list proved effective in decreasing the presence of non-authoritative sources on English Wikipedia, later research revealed significant limitations when applying this concept to other language editions (Baigutanova et al., 2023b). A few other language communities have developed their own versions of the perennial source list, but such cases are rare among the hundreds of Wikipedia editions, and these lists typically contain only limited data. Furthermore, sources flagged as unreliable in one language edition are often still used in others, highlighting the variability in how source reliability is perceived across cultural and linguistic contexts. These discrepancies pose a major challenge to the development of universally applicable models for assessing source reliability across Wikipedia language editions.
@@
The multilingual nature of Wikipedia calls for specific modeling techniques (Johnson and Lescak, 2022). Models such as multilingual BERT have been used to detect vandalism (Trokhymovych and Saez-Trumper, 2021) or political bias (Swati et al., 2023), while other works have built rankings of relative quality in multiple language versions of the platform (Lewoniewski et al., 2020, 2019). To build a multilingual model, such as one recently proposed for entity insertion (Feith et al., 2024), one needs a large dataset of training data in all the languages of interest. For some other tasks, and due to the limitations found in languages with smaller Wikipedia datasets, language-agnostic approaches. These approaches have proven effective across a wide range of tasks, e.g., the assessment of article content quality (Das et al., 2024), reliability (Wong et al., 2021), automatic entity-linking (Gerlach et al., 2021), orphan article detection (Arora et al., 2024), and topic modeling (Piccardi and West, 2021) and classification (Johnson et al., 2021).
Language-agnostic models using features extracted from the network structure or user behavior have also been used in the more general context of detecting misinformation or low-quality information in social media. In this line, Shu et al. (2019) constructed embeddings of users and news and used diffusion models to trace the paths through which fake news propagates, while Zhao et al. (2021) extracted features from the user interaction network to help identify misinformation in online health communities. As shown later, user-behavior related language-agnostic signals are useful for modeling source reliability, and can be adapted across languages to improve performance in low-resource settings.
@@ 3.1. Perennial Sources
We use the English Wikipedia perennial source list (Wikipedia contributors, 2024) as ground truth for modeling source reliability. As only a few other languages have perennial source lists, in this work we focus on the list from English Wikipedia, as it is the most complete one, and concerns the largest language Wikipedia. We discuss this lack of ground truth data in the Limitations section below. It was collected on March 30, 2023. Although the list exists for 12 other languages, we only use the English version as it is by far the most extensive (see (Baigutanova et al., 2023b) for details on coverage). In English, it consists of a collection of web domains in five categories: blacklisted (569), deprecated (125), generally unreliable (178), no consensus (137), and generally reliable (184). To build a binary classifier, we use the generally reliable domains as positive labels, merge the domains blacklisted, deprecated, and generally unreliable into a single unreliable category (i.e., the negative label), and ignore the no consensus domains.
Table 1. Statistics of the topic-language datasets used in the study grouped by high, mid, and low resource languages. All numbers are averages (apart from the number of languages). Perennial Domains (P. Domains) refers to the number of domains for the perennial source list in English found in the different datasets. See Table A1 in Appendix B for the corresponding median values.
Topic Langs Articles Revs Revs/Ar URLs URLs/Ar Domains P. Domains
High Climate Change 7 $1378$ 372 133372\,133 $269$ 58 67158\,671 $49$ 15 36315\,363 $203$
COVID-19 7 $898$ 233 905233\,905 $260$ 60 16860\,168 $89$ $8289$ $206$
Biology Sample 7 13 60713\,607 586 171586\,171 $43$ 59 76359\,763 66 11 71311\,713 $148$
History Sample 7 $5710$ 624 428624\,428 $109$ 52 70252\,702 $12$ 12 12112\,121 $166$
Media Sample 7 $5577$ 819 541819\,541 $146$ 108 583108\,583 $22$ 21 29021\,290 $259$
Mid Climate Change 36 $395$ 28 36428\,364 $71$ $6224$ 1919 $2599$ $94$
COVID-19 37 $272$ 18 52318\,523 $68$ $8984$ $39$ $1788$ $102$
Biology 37 $4153$ 83 59983\,599 2020 $8691$ 33 $2022$ $61$
History 37 $1680$ 83 25883\,258 $49$ $6274$ 77 $2054$ $67$
Media 37 $1018$ 46 95746\,957 $46$ $7932$ $10$ $2416$ $123$
@@
Biology 41 $593$ 21 09621\,096 $35$ $1159$ 44 $460$ $16$
History 55 $511$ 13 50113\,501 $26$ $952$ 66 $406$ $22$
Media 70 $159$ $2262$ $14$ $528$ 77 $264$ $30$
@@ 3.2. Language Groups
As we aim to test the proposed model on a variety of languages, we collect data on all $326$ available language editions of Wikipedia. We then discard all languages that have fewer than two reliable and unreliable sources according to the English perennial source list. Among the remaining ones, we distinguish between low, middle, and high-resource languages by examining the number of “active users”, defined as “registered users who have made at least one edit in the last $30$ days.” [^4] https://meta.wikimedia.org/wiki/List_of_Wikipedias (accessed on 01-03-2024) We consider the top 5% of the languages having the most active users as high-resource, the subsequent 25% as mid-resource, and the remaining 70% as low-resource languages.
@@ 3.3.1. Topic Selection
In this study, we have focused on articles from the following five topic datasets: Climate Change, COVID-19, Biology, History, and Media. Climate Change and COVID-19 were selected as they are both controversial, but for different time durations. COVID-19 appeared at the end of 2019, while Climate Change has been a controversial topic for decades (Korte et al., 2023). Both have a Wikiproject community of editors  [^5] https://en.wikipedia.org/wiki/Wikipedia:WikiProject who label articles as relevant to the topic with special guidelines. For comparison, we also selected three topics that are not as controversial, but which varied in terms of reference quality based on the “reference risk” metric. This metric was introduced by Baigutanova et al. (2023a) and estimates the likelihood of articles citing unreliable sources. Specifically, reference risk is calculated as the proportion of citations in an article that point to domains categorized as generally unreliable, deprecated, or blacklisted in the English Wikipedia perennial source list. Higher reference risk indicates a greater potential threat to verifiability and information quality.
This stratification allowed us to evaluate our model’s performance across topics with varying levels of citation quality and editorial scrutiny.
@@ 3.3.2. Data Download
For the first two topics, we use the articles listed by the corresponding topic-based WikiProjects. For the latter three, we use the topics identified using the Wikipedia Objective Revision Evaluation Service (ORES) machine learning system (Halfaker and Geiger, 2020). We use this method rather than later approaches (Johnson et al., 2021), since it is the one employed in Baigutanova et al. (2023a), which inspires our comparison among topics. To keep the sizes of datasets comparable and manageable, we sample the articles identified by ORES to be in Biology and History topics by 10% (each), and in the Media topic by 2%. These two approaches allow us to find articles in the English Wikipedia for each of the five topics, resulting in a total of 84 02784\,027 articles. To find the versions of these articles in other languages, we use the MediaWiki API, which results in another 335 809335\,809 of articles in all of the 326 languages. We collect the articles and their complete edit history using the MediaWiki API in between December 2023 and February 2024.
@@ 3.3.3. Extraction of Source Domains
Citations may appear in the body of the article text and as references at the bottom of the article, each of which has a standard template. Each revision that involves a citation contains the author of the revision, position within the document of the citation, text of the citation, and meta information, such as whether it is a DOI, URL, ISBN, ISSN, etc. or a raw URL. We standardize these revisions to create a list of “source edits” containing the metadata about each insertion, deletion, or edit that involves a source. We resolve redirects, apply rules to clean and standardize the URLs (including DOI references redirecting to a resource), and merge revisions when the same user edits an article in succession (regardless of the time). Finally, we extract the domains from the URLs, obtaining 376 566376\,566 unique source domains.
@@ 3.3.4. Dataset Statistics
The summary statistics about the topic-language datasets (or simply “datasets” throughout this paper), for the five topics grouped into low-, mid-, and high-resource languages can be seen in Table 1. The medians of these values can be found in Appendix B, while more in-depth statistics for each individual language are in Appendix C. Recall that we only use datasets that have passed the threshold of having at least 2 reliable and 2 unreliable perennial sources.
@@ 3.4. Feature Definition
Unlike previous studies measuring Wikipedia edit quality that did not consider certain types of edits and reverts (Trokhymovych et al., 2023), we propose to focus on the entire edit history concerning each domain. Using edit metadata, we define 52 features capturing the popularity of the domain, the “permanence” of the domain across edits, and the number and type of users involved in edits adding or removing that domain. As the features are computed for each domain within each dataset separately (recall, by “dataset” we mean a collection of articles on a certain topic, in a certain language) and the datasets are of varying sizes and editorial frequencies, we explore different ways of normalizing these statistics w.r.t. the first appearance of the domain, age of the dataset, and in terms of time duration or number of revisions. Below, we give a short overview of the features we have used. A more detailed description can be found in Appendix A.